### v1

# GRiNS: A Python Library for Simulating Gene Regulatory Network Dynamics

Abstract
The emergent dynamics of complex gene regulatory networks govern various cellular processes. However, understanding these dynamics is challenging due to the difficulty of parameterizing the computational models for these networks, especially as the network size increases. Here, we introduce a simulation library, Gene Regulatory Interaction Network Simulator (GRiNS), to address these challenges. GRiNS integrates popular parameter-agnostic simulation frameworks, RACIPE and Boolean Ising formalism, into a single Python library capable of leveraging GPU acceleration for efficient and scalable simulations. GRiNS extends the ordinary differential equations (ODE) based RACIPE framework with a more modular design, allowing users to choose parameters, initial conditions, and time-series outputs for greater customisability and accuracy in simulations. For large networks, where ODE-based simulation formalisms do not scale well, GRiNS implements Boolean Ising formalism, providing a simplified, coarse-grained alternative, significantly reducing the computational cost while capturing key dynamical behaviours of large regulatory networks. The documentation and installation instructions for GRiNS can be found at https://moltenecdysone09.github.io/GRiNS/.
Keywords:
Gene Regulatory Networks, Network Dynamics, Parameter-Agnostic Simulation, Ising Boolean Formalism, Random Circuit Perturbation (RACIPE), Systems Biology
1 Introduction
One of the core goals of systems biology is to understand and model complex interactions in a biological system to understand the mechanistic underpinnings of various cellular processes. The advent of high-throughput technologies capable of capturing the molecular profiles of cells has enabled us to infer causal relations between molecules governing various biological processes. A class of such models that deal with the mechanistic relation between the genes that govern a cellular process are gene regulatory networks (GRNs). GRNs are network representations of the causal interactions between genes that govern their expression levels. GRNs are crucial for understanding the mechanistic details governing complex biological processes that give rise to specific expression patterns and cellular phenotypes [1, 2, 3, 4]. As representations of the interactions between genes, GRNs provide a framework to decipher the regulatory logic and identify key interactions that determine a particular biological outcome. There are two main approaches to constructing GRNs. The bottom-up approach employs detailed experiments to verify the interactions between genes to build these GRNs, with experiments to tease out causal relations between genes. However, while being precise, this approach is cumbersome because, with the increasing network size, it becomes complicated to accurately determine all the interactions between the genes. On the other hand, the top-down approach leverages high-throughput technologies and the vast amounts of data they generate to infer the interactions between genes. Various algorithms have been developed that utilize and integrate data from different single-cell molecular profiling techniques to infer gene regulatory networks [5, 6, 7, 8]. However, unlike the bottom-up approach, such inferred GRNs often have limited predictive potential due to missing or inaccurate regulatory links [9].
When it comes to understanding the dynamics of such predicted GRNs, most of these methods suffer from a common challenge of being unable to accurately predict the interaction functions and their parameters [9, 10]. Given the uncertainty in data due to biological noise and technical limitations and considering the sizes of the large networks that govern these interactions, it becomes challenging to parameterize these networks accurately. Having simulation frameworks that are parameter agnostic and focus on the dynamics and steady states of a given network structure becomes important in such cases. These approaches do not depend on specific parameter sets to explain the behavior of GRNs. An advantage of such approaches is that, even in the absence of parameters, having qualitative estimates of the general ranges of the parameters is good enough for getting an idea about the dynamics of the network in the parameter range.
RAndom CIrcuit PErturbation (RACIPE) is one such framework that tries to identify the possible phenotypic space of a network given its topology [11, 12]. It samples parameters over predefined ranges and simulates them over multiple initial conditions. The tool can capture a network’s steady states by randomly sampling the parameters and simulating them over multiple initial conditions. This approach helps one to understand the possible types of dynamics and steady states a network can show, even when the mathematical model’s precise kinetic parameters are absent. Additionally, due to the random sampling of parameters and initial conditions, the RACIPE tool mimics the wide range of variability observed in biology, giving us a more realistic view of the network’s behavior compared to when we use precise parameters, which, more often than not, are very context-specific and may not represent the actual behavior of the network in a different scenario. As RACIPE automates the entire pipeline of model building and its subsequent simulation according to the inputs provided by the user, it proves to be a suitable tool for analyzing the dynamics of GRNs.
Another such parameter-agnostic method is the Ising boolean simulation framework [13]. This framework represents each gene as a binary variable—active or inactive— depending on the cumulative influence of the incoming active links. Since each update step is based on matrix multiplication, this method, although very crude, is suitable for simulating large networks where the system of ODEs is too large and, hence, is too slow to be practical to be simulated over thousands of parameters and initial conditions through RACIPE.
Since both these methods depend heavily on matrix-based operations, they are ideal candidates for GPU implementation. GPUs are suitable for matrix-based operations because they can parallelize the computation, giving dramatic speed-ups compared to the same computations run on a CPU. The Python library Jax’s efficient array-oriented numerical computation functions provide an excellent foundation for our simulation library [14]. RACIPE, in particular, requires GPU-based differential equation solver implementations provided by the diffrax [15] library, which is built on top of Jax. Utilizing these, we introduce a simulation library for GRNs that exploits the speed of GPUs over CPUs for matrix operations, enabling rapid and scalable simulations of GRNs. Our library’s modular design, written in an easily accessible language like Python, offers more options for the users to tweak the simulations according to their needs.
Figure 1: Overview of the simulation frameworks in GRiNS. GRiNS includes implementations of Random Circuit Perturbation (RACIPE) for continuous ODE-based modeling and Ising Boolean formalism for discrete-state simulations.
2 Methods
2.1 RAndom CIrcuit PErturbation (RACIPE)
RACIPE is a GRN modeling framework designed to sample the steady-state repertoire of a GRN purely based on its topological structure [11, 12]. It does this by generating a system of coupled ordinary differential equations (ODEs) to represent the interactions of the genes and then simulating these equations over multiple randomly sampled parameter sets and initial conditions. By simulating a parameterized system of ODE over the initial conditions, RACIPE can uncover the possible set of steady states. When this process is carried out on all the parameter sets, RACIPE can map out the possible phenotypic outcomes of a network even in the absence of specific parameter sets.
In the following sections, we describe the methodology of RACIPE and how it constructs the differential equation models from network topologies and systematically samples parameters within biologically relevant ranges. We will then describe the simulation methodology of the parameterized ODEs over the initial conditions and the subsequent analysis, which can be done to identify robust, steady states to uncover the dynamic landscape of gene regulatory networks.
2.1.1 Parsing GRNs to Construct System of ODEs
RACIPE models a signed and directed GRN as a system of coupled ODEs. For a gene, T, in the GRN, the ODE describing the change in its expression value as a function of the input nodes is given by:
$\begin{split}\frac{dT}{dt}=G_{T}*\prod_{i}\frac{H^{S}(P_{i},{P_{i}}^{0}_{T},n_{P_{i}T},\lambda_{P_{i}T})}{\lambda_{P_{i},T}}*\prod_{j}H^{S}(N_{j},{N_{i}}^{0}_{T},n_{N_{j}T},\lambda_{N_{j}T})-k_{T}*T\end{split}$ (1)
Where,
$\begin{split}H^{S}(B,B^{0}_{A},n_{BA},\lambda_{BA})=\frac{{B^{0}_{A}}^{n_{B,A}}}{{B^{0}_{A}}^{n_{BA}}+B^{n_{BA}}}+\lambda_{BA}*\frac{B^{n_{BA}}}{{B^{0}_{A}}^{n_{BA}}+B^{n_{BA}}}\end{split}$ (2)
$G_{T}$ in Equation (1) refers to the maximal expression value of the node T, $H^{s}$ is a modified form of the hill’s equation called the shifted hills equation, as given in the Equation (2). Each shifted hills equation (2) represents the effect of an upstream incoming node on T. It has the threshold parameter $B^{0}$, $n_{BA}$ is the hill’s coefficient, and $\lambda_{BA}$ is the fold change parameter representing the fold change in the expression of A brought about by the effect of the edge from node B. $P_{i}$, and $N_{j}$ refers to the values of the input activating and inhibiting nodes respectively. All the hill’s terms of the incoming edges are multiplied with the maximal expression value $G_{T}$ to get a scaled value of T’s expression, the production term of the ODE. Additionally, $k_{T}$ is the degradation rate of the node and is multiplied by the value of T to get the degradation term of the equation.
The current parser supports only signed and directed GRNs, i.e., it only recognizes activation and inhibition links when constructing the system of ODEs. Once generated, the ODE system is written into a Python file formatted for compatibility with the Diffrax simulation library. Users can modify these functions to incorporate custom dynamics, as the subsequent simulation steps do not depend on the original RACIPE ODE structure given above.
2.1.2 Sampling Parameters and Initial Conditions
Since RACIPE’s primary object is to capture all possible steady states of a given GRN, the parameter sampling strategy must eliminate any biases that may skew the result. The number of parameters to be sampled for a network with N nodes and E edges is 2N + 3E. The term 2N represents the maximal production and degradation rate parameters that must be sampled. The 3E term represents the threshold, the hill’s coefficient, and the fold change parameters that need to be sampled for each of the edges present in the network. Table 1 gives the default ranges of the parameter values over which the sampling is done. Th
Parameters Minimum Maximum
Production Rate (G) 1 100
Degradation Rate (k) 0.1 1
Fold Change (Inhibition $\lambda$) 0.01 1
Fold Change (Activation $\lambda$) 1 100
Hill Coefficient (n) 1 6
Threshold The ranges depend on the in-degree - half functional rule
Table 1: Default parameter ranges used by RACIPE.
The inhibition fold change parameter is sampled in the inverse range, i.e., between the inverse of the maximum value and the inverse of the minimum value. Following this, the inverse of all the sampled values is taken. This process shifts the mean of the distribution from 0.5 to 0.02 (in the case of default ranges), which ensures that both weak and strong inhibitory fold change parameters are sampled in a balanced manner.
Another aspect to consider to eliminate biased parameters of nodes is the threshold values assigned to the edges. For nodes with incoming edges, the threshold values need to be sampled to be within the minimal and maximal steady-state expression ranges of the nodes from which they originate. Otherwise, a deviation from this would mean that an edge could be always active or always inactive, biasing the simulations. The half-functional rule is employed to correct this bias, and it ensures that the threshold values are chosen between 0.02 and 1.98 times the median steady-state expression level of the upstream node. Suppose other genes regulate the upstream node itself. In that case, its steady-state expression distribution is determined by simulating its steady-state levels based on the regulatory parameters of its incoming nodes. The median expression of the upstream node is then used to define the appropriate range for threshold sampling. This ensures that the threshold values are chosen so that each regulator has a probability of being active at roughly 50% across all the simulations.
The values are sampled from the respective node’s minimum and maximum expression values for the initial conditions. Our framework supports multiple sampling methods—including Sobol (default), Latin hypercube, uniform, log-uniform, normal, and log-normal distributions—and even allows mixing different distributions for different parameters or initial conditions [16]. This flexibility lets users target specific regions of parameters or the initial condition space. A parameter range file is also generated for easy customization, on which the parameters can be regenerated for subsequent simulations. Similarly, initial conditions can also be regenerated based on the updated ranges.
2.1.3 Simulating the ODE system and processing the results
Once the parameters and initial conditions are generated after the GRN is parsed to create the ODE file, the library provides functions for the simulation of the ODE system for all combinations of parameter sets and initial condition values. A function is also provided to simulate the ODE system for a single combination of parameters and initial conditions. The simulations are usually done using the vmap function of the Jax library, and the user can control the batch size for cases where the VRAM is insufficient to run large-scale simulations. The user can control the tolerances (relative and absolute), the start time point (by default is zero as we would be dealing with initial conditions), and the end time point of the simulations, after which the simulation will terminate even though the steady-state may not have been reached. By default, the simulations terminate once the steady state condition is reached, which is determined by the tolerance values. The resulting data—comprising node values, termination times, a steady-state indicator, and identifiers for each parameter–initial condition pair—is stored in a data frame as a Parquet file on the disk.
Setting the end time point sufficiently high is important to allow most combinations to reach a steady state, which the steady state flag column of the solution data frame can track. The implementation also supports recording gene expression levels at custom time points for time-series analyses. However, we recommend first determining the steady-state range and then deciding on the optimal spacing of the time points and their range to capture the relevant dynamics.
Since parameters can vary across each simulation instance, a normalization method is necessary to compare steady-state values. Our library provides functions that normalize node expression by the maximal expression-to-degradation ratio (G/k) – the highest possible expression level – resulting in outputs scaled between 0 and 1, with 1 being the maximal expression of the node. Additionally, nodes can be discretized into levels based on the global distribution of normalized G/k values, offering a standardized approach to processing simulation outputs. These functions are compatible with both the steady-state and time-series simulation results.
2.2 Ising Formalism
Although RACIPE formalism is an excellent framework for understanding the possible dynamics emergent from a GRN, computational cost becomes prohibitive for large networks due to the increasing number of parameters and longer simulation times. Ising formalism provides a simpler, albeit coarse-grained approach for simulating large networks [13]. In this approach, the state of the variable is represented by a discrete variable indicating the on or off state of the gene. The system starts with an initial condition ($s_{0}$), and the subsequent state ($s_{t+1}$) is obtained by the previous state’s ($s_{t}$) matrix product with the network’s adjacency matrix (A), where -1 represents inhibitory interactions, 1 represents activating interactions, and 0 indicates no connection. Once the state is multiplied with the adjacency matrix, the resultant vector is then converted to 1s and -1s (or 0s depending on the flip values provided) according to the rule given below:
$s_{i}(t+1)=\begin{cases}1,&\sum\limits_{j}J_{ij}s_{j}(t)>0\\ s_{i}(t),&\sum\limits_{j}J_{ij}s_{j}(t)=0\\ -1,&\sum\limits_{j}J_{ij}s_{j}(t)<0\end{cases}$ (3)
The system can be updated using either a synchronous update, where all the nodes are updated simultaneously, or an asynchronous update, where one randomly chosen node is updated at each step. The user can set the number of simulation steps, define initial conditions (or use randomly generated ones by default), and specify a custom update order for asynchronous mode. The simulation results are stored in a dataframe with columns for the binary node states (0 or 1), the simulation step number, and the corresponding initial condition number. Optionally, the node values at each step can be stored as bytes to conserve disk space and later unpacked during analysis.
3 Conclusion
GRiNS uses parameter-agnostic simulation frameworks and sampling strategies to model GRNs and capture their dynamics under various conditions. Implementing these methods in an accessible language like Python coupled with modular functions allows users to customize their simulations and explore the dynamic properties of GRNs in a standardized and reproducible manner. Thanks to its GPU-based simulation, GRiNS can be easily integrated into machine learning pipelines for cases like network inference, where unbiased simulation methods are important to determine and evaluate the predicted networks. The documentation and installation instructions for GRiNS can be found at https://moltenecdysone09.github.io/GRiNS/.
Competing Interests
The authors declare no competing interests.
Author Contributions Statement
M.K.J. and P.H. conceived the idea of the simulation package. P.H. and H.B.V. wrote and developed the package. M.K.J. and P.H. wrote and reviewed the manuscript
Acknowledgments
M.K.J. received support from Param Hansa Philanthropies. P.H. and H.B.V. were supported by their respective Prime Minister’s Research Fellowships (PMRF) awarded by the Government of India.